HETAV PATEL

Grade 11,

Find an AP whose first term is unity and the second , tenth and 34th term form a GP

Find an AP whose first term is unity and the second , tenth and 34th term form a GP

Grade:11

1 Answers

Ravi
9 years ago
For AP
a=1
2nd term(AP) = a+d =1st term of (GP)= b1
10th term(AP) = a+9d =2nd term of (GP)=b2
34th term(AP) = a+33d =3rd term of (GP)=b3


For a GP

(b2)^2 = b1*b3
Put the values, form a quadratic equation, put a=1 and dfind the value of d.